OpenStudy (anonymous):

Estimate the area under the curve cos sqrt x for 0=

OpenStudy (anonymous):

Definite Integrals? Any help..im stuck

OpenStudy (anonymous):

Use trapezoidal rule. If you look at a plot of your function, you'll see that it falls roughly linearly from 0 to 2.

OpenStudy (anonymous):

The answer with trap. rule would be about 1.156

OpenStudy (anonymous):

the problem is we havnt got to trapeziod rule

OpenStudy (anonymous):

Not sure why it's asking you to estimate when you can solve it exactly with a u sub then by parts.

OpenStudy (anonymous):

\[\int\limits_{a}^{b} f(x)dx \approx (b-a)\frac{f(a)+f(b)}{2}\]

OpenStudy (anonymous):

Well then you have to estimate with rectangles.

OpenStudy (anonymous):

You should know that the area of a trapezoid is \[\frac{1}{2}(a+b)h\]where a and b are the 'top' and 'bottom' and h is the altitude (or height). Here, a=f(0), b=f(2) and the height is (b-a)=2

OpenStudy (anonymous):

It's okay to apply this formula.

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

np